Problem to solve is: z^1/2 - 4z^1/4 + 4 = 0
so I let u=z^1/4 and I let u^2=z^1/2
Then I get u^2-4u+4=0 which factors to (u-2)^2
So U-2=2
U=2
My question is how do I check this. I am confused. May I please ask for help?
Thank you

Answer by scott8148(6628)

2 = z^(1/4) ___ 16 = z

16^(1/2) - 4[16^(1/4)] + 4 = 0 ___ does this check?